IQs are normally distributed with mean of 100 and a standard deviation of 15.

Fill in the blank

85% of the population have an IQ that is greater than _______? (2 decimals)
Answer #1

solution

Using standard normal table,

P(Z > z) = 85%

= 1 - P(Z < z) = 0.85

= P(Z < z ) = 1 - 0.85

= P(Z < z ) = 0.15

z = -1.04 (using standard normal (Z) table )

Using z-score formula  

x = z * +

x= -1.04 *15+100

x= 84.40